OpenStudy (anonymous):

3/4j + 1 great than 4 3/4j +1 -1 great than 4 -1 3/4j great than 3 4/3(3/4)j great than 3(4/3) j great than 4

OpenStudy (anonymous):

Looks right to me. As long as the original problem is \[(3/4)j+1 \ge 4\]

OpenStudy (anonymous):

no it is only great than not great than or equal to where did I go wrong?

OpenStudy (anonymous):

My bad, you're right. i put the wrong sign

OpenStudy (anonymous):

ok yeah

OpenStudy (anonymous):

Do you have to put this one on the number line?

OpenStudy (anonymous):

yes I was just going to ask it would be a close circle on 4 and shaded to the left

OpenStudy (anonymous):

You use closed circle when its greater than or EQUAL to use open circle when its just greater than/less than and greater than means shading to the right

OpenStudy (anonymous):

I am sorry I have another one do you mind?

OpenStudy (anonymous):

that makes sence now thanks

OpenStudy (anonymous):

oops sense

OpenStudy (anonymous):

- 5/6x +3 great than 1/2 -5/6x +3-3 great than 1/2(-3) ok I am stuck

OpenStudy (anonymous):

on 1/2 - 3?

OpenStudy (anonymous):

yes I am drawing a blank is it -3/2

OpenStudy (anonymous):

put -3 over 1 (1/2) - (3/1) now find a common denominator which is 2 (1/2) - (6/2) subtract straight across the top and keep the denominator

OpenStudy (anonymous):

I just made that hard

OpenStudy (anonymous):

1-6=-5 so it is -5/2 so now you have (-5/6)x > -5/2

OpenStudy (anonymous):

so than it would be -5/6x great than -5/2 6/5(-5/6) x great than -5/2(6/5) x great than -10

OpenStudy (anonymous):

you divided by a negative, so it flips the sign x<(-5/2)*(-6/5)

OpenStudy (anonymous):

oh ok otherwise I am good?

OpenStudy (anonymous):

No. (-5/2)*(-6/5)=3 so x<3

OpenStudy (anonymous):

do we change the sign back because it became positive or still leave it flipped?

OpenStudy (anonymous):

You leave it. we divided by -5/6 so it flipped.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I have a couple moew just want to make sure I am doing them right do you mind?

OpenStudy (anonymous):

2(b-7) + -4b greater than 30 -18 2b - 14 + -4b great than 12 -14 -2b great than 10 14 -14 -2b great than -4 b less than -8

OpenStudy (anonymous):

sorry I think I messed up the last part 14-14-2b great than 24 b less than 12